Design a 35-km pipe (e=.00015m) to carry petroleum (sg = 0.82) with head losses of 5.4m per 1000m. viscosity = 1.2x10^-6 m^2/s. Q = 0.16cms
Answer #1


Given headloss she. = 5-4/10oom. s Total headlos în 35kn 5 he = 5.4*15= 189m by darcys lan 189 = fx 35000 x 10.16)2 ? 12.1X So check the = 0.82 *0.16*0.3454 403032.78710 /4X(0.14592 х1-2010 b =0.345 not correl. for Greater, Re in Smooth pipe VF V 2.